solve for x: log8 (x+1)=2/3
Answer by lwsshak3(11628)

solve for x: log8 (x+1)=2/3
..
log8 (x+1)=2/3
convert to exponential form
8^2/3=x+1
take the cube root of 8=2,then square it=4
4=x+1
x=3